OpenStudy (ryguy):

How do you find the solution set of (x - 3) - 2(x + 6) = -5

OpenStudy (sleepyhead314):

first distribute the (-2) into the (x+6) can you do that? :)

OpenStudy (ryguy):

yeah -2x + -12

OpenStudy (sleepyhead314):

good! :) so now we have x - 3 - 2x - 12 = -5 can you add like terms?

OpenStudy (ryguy):

(-12)*

OpenStudy (anonymous):

no x=-10

OpenStudy (ryguy):

3x -15= -5

OpenStudy (ryguy):

skate go away im not trynna cheat. im trying to figure it out so i can do it without cheating

OpenStudy (sleepyhead314):

oooo x - 2x = ? careful ^_^

OpenStudy (anonymous):

https://mathway.com/

OpenStudy (ryguy):

oh so does finding the solution set basically just mean solve the equation? the solution set should be {-10} correct?

OpenStudy (sleepyhead314):

yes ^

OpenStudy (ryguy):

cool thanks sleephead :D

OpenStudy (sleepyhead314):

you're welcome ^_^